# Closed-Form Feedback-Free Learning with Forward Projection

### v1

@@ A.3 Implementation of Local Supervision and Forward Forward
The Forward-Forward algorithm was implemented by generating both “positive” data-label pairs $[x,y]$ and “negative” data pairs $[x,\check{y}]$ by concatenation. Positive activations were generated such that
@@
In MLP implementations, concatenation was implemented with supplementary input neurons to hold the label information as a one-hot vector. In convolutional implementations, the label was concatenated in the channel dimension, such that each channel indicated a single class, with constant value over all kernel positions.
A.4 Analysis of label projection approaches
@@
$\textbf{W}_{l}=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}+\lambda I)^{-1}\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}.$ (13)
We let $\operatorname{rank}(\tilde{\textbf{Z}})$ denote the column rank of $\tilde{\textbf{Z}}$ – the maximum number of linearly independent column vectors in its column space. Since $\textbf{W}_{l}$ is a linear function of $\tilde{\textbf{Z}}$, we have $\operatorname{rank}(\textbf{W}_{l})\leq\operatorname{rank}(\tilde{\textbf{Z}})$.
@@
A.5 Computing the Forward Projection estimator
@@
$\textbf{W}_{l}=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}+\lambda I)^{-1}(\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}).$ (18)
A.6 Information Encoding and Interpretability
@@
A.7 Modelling with Challenging Hidden Activation Functions
@@
Table 3: Test accuracies of local learning methods in FashionMNIST, Promoters and PTBXL-MI tasks, using modulo 2 (“mod2”) ($f(x)=x\mod{2}$) and square ($f(x)=x^{2}$) activation functions. FP: Forward Projection; LS: Local Supervision; FF: Forward-Forward; BP: Backpropagation
A.8 Few-shot modelling
@@
Performance of forward projection, random features, local supervision, forward-forward and backpropagation in few-shot modelling tasks is provided in Supplementary Table 4.
Dataset Partition Training Samples
CXR Train N=5 100.0 ± 0.0 100.0 ± 0.0 50.7 ± 16.0 52.0 ± 9.5 72.5 ± 9.4
N=10 100.0 ± 0.0 100.0 ± 0.0 48.7 ± 12.7 48.8 ± 6.3 69.0 ± 8.1
N=15 100.0 ± 0.0 100.0 ± 0.0 55.2 ± 13.9 51.1 ± 6.1 72.4 ± 10.6
N=20 100.0 ± 0.0 100.0 ± 0.2 58.5 ± 12.8 50.0 ± 5.5 72.6 ± 8.0
N=30 100.0 ± 0.0 99.9 ± 0.4 59.8 ± 15.5 50.0 ± 5.7 71.8 ± 9.6
N=40 100.0 ± 0.1 99.3 ± 0.8 61.4 ± 15.8 49.7 ± 4.8 74.3 ± 10.2
N=50 99.9 ± 0.2 98.4 ± 1.0 66.1 ± 14.8 49.4 ± 6.2 84.2 ± 9.8
Test N=5 75.2 ± 7.0 67.6 ± 7.5 49.7 ± 8.8 49.3 ± 5.7 57.6 ± 7.6
N=10 76.6 ± 5.8 68.1 ± 6.0 47.9 ± 9.0 49.1 ± 5.7 59.7 ± 6.2
N=15 77.4 ± 4.5 69.2 ± 4.8 53.9 ± 12.3 49.9 ± 6.7 63.1 ± 8.6
N=20 78.9 ± 3.7 69.0 ± 4.9 58.9 ± 11.5 49.0 ± 5.8 66.1 ± 9.3
N=30 80.6 ± 3.7 69.5 ± 3.7 59.2 ± 14.0 51.2 ± 6.2 66.6 ± 8.0
N=40 81.7 ± 2.9 72.4 ± 3.7 61.2 ± 16.1 48.0 ± 6.4 70.2 ± 8.3
N=50 82.5 ± 3.5 72.2 ± 3.1 64.8 ± 14.9 49.4 ± 7.4 78.9 ± 8.3
OCT Train N=5 100.0 ± 0.0 100.0 ± 0.0 53.9 ± 27.2 49.4 ± 12.5 86.8 ± 13.2
N=10 100.0 ± 0.0 100.0 ± 0.0 58.3 ± 17.8 50.4 ± 12.2 83.3 ± 12.2
N=15 100.0 ± 0.0 100.0 ± 0.0 57.1 ± 15.7 49.9 ± 11.2 80.5 ± 11.2
N=20 100.0 ± 0.0 100.0 ± 0.0 53.4 ± 21.0 51.6 ± 9.8 88.9 ± 12.1
N=30 100.0 ± 0.0 100.0 ± 0.0 58.0 ± 18.8 50.1 ± 11.3 83.7 ± 12.3
N=40 100.0 ± 0.0 99.9 ± 0.3 61.4 ± 21.0 52.3 ± 9.2 88.0 ± 12.0
N=50 100.0 ± 0.0 99.7 ± 0.7 61.7 ± 19.8 50.8 ± 13.6 95.2 ± 8.3
Test N=5 76.3 ± 11.5 63.1 ± 10.0 51.1 ± 17.6 52.5 ± 12.0 61.8 ± 17.4
N=10 84.5 ± 5.8 66.5 ± 7.7 54.6 ± 16.7 51.8 ± 12.1 71.8 ± 15.9
N=15 86.5 ± 7.8 68.7 ± 7.4 56.8 ± 16.3 47.8 ± 12.8 72.3 ± 14.4
N=20 88.4 ± 5.9 69.5 ± 6.0 49.6 ± 19.1 52.3 ± 13.0 82.4 ± 14.4
N=30 89.9 ± 4.0 70.5 ± 5.7 55.7 ± 20.9 51.8 ± 11.7 81.9 ± 12.1
N=40 90.2 ± 5.0 73.3 ± 5.5 62.0 ± 25.1 50.6 ± 13.0 88.7 ± 9.3
N=50 91.2 ± 4.7 73.6 ± 5.9 62.6 ± 23.8 49.7 ± 18.4 94.2 ± 6.6
@@
Table 4: Train and Test AUC of local learning methods and backpropagation in few-shot learning experiments. CXR: chest x-ray dataset. OCT: optical coherence tomography dataset.

### v2

@@ A.4 Implementation of local learning methods
The Forward-Forward algorithm was implemented by generating both “positive” data-label pairs $[x,y]$ and “negative” data pairs $x,\check{y}]$ by concatenation. Positive activations were generated such that
@@
In MLP implementations, concatenation was implemented with supplementary input neurons to hold the label information as a one-hot vector. In convolutional implementations, the label was concatenated in the channel dimension, such that each channel indicated a single class, with a constant value over all kernel positions. Predictive Coding [8, 5] was implemented using inverse layers $\textbf{W}_{l}^{\operatorname{Back}}\in\mathbb{R}^{m_{l}\times m_{l-1}}$ for each weight matrix $\textbf{W}_{l}$, to estimate the pre-synaptic inputs from the post-synaptic outputs, such that
$\hat{\textbf{a}}_{l-1}:=f_{l}^{\operatorname{Back}}\left(\textbf{a}_{l}\textbf{W}_{l}^{\operatorname{Back}}\right)$ (17)
The local loss is given by the reconstruction error
$\mathcal{L}_{l}:=\lVert\textbf{a}_{l-1}-\hat{\textbf{a}}_{l-1}\rVert_{2}^{2}$ (18)
Similarly, Difference Target Propagation [1] aims to reconstruct presynaptic inputs, such that
$\hat{\textbf{a}}_{l-1}:=\textbf{a}_{l-1}-f_{l}^{\operatorname{Back}}\left(\textbf{a}_{l}\textbf{W}_{l}^{\operatorname{Back}}\right)+f_{l}^{\operatorname{Back}}\left(\hat{\textbf{a}}_{l}\textbf{W}_{l}^{\operatorname{Back}}\right)$ (19)
In our implementations, we set $f_{l}^{\operatorname{Back}}=\operatorname{ReLU}$ in both Predictive Coding and Difference Target Propagation models.
A.5 Analysis of label projection approaches
@@
$\textbf{W}_{l}=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}+\lambda\textbf{I})^{-1}\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}$. (20)
[…]
@@
A.6 Computing the Forward Projection estimator
@@
$\textbf{W}_{l}=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}+\lambda\textbf{I})^{-1}(\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}).$ (25)
A.7 Information Encoding and Interpretability
@@
A.8 Stability Experiments
Technical analysis and stability experiments were performed on MLP models training on the FMNIST data. The stability of layer explanations, on models fitted to different data folds, with different random projection matrices, using a $4\times 1000$ neuron feedforward architecture was evaluated. Test performance of layer explanations improved progressively from the first layer to the final layer with low variability throughout (Figure 6-A). Model stability with respect to choices of the FP penalty parameter $\lambda$ (Equation 2) in hidden layers was evaluated in a $3\times 1000$ neuron MLP. Test accuracies were stable from $\lambda=1.25$ to $\lambda=80$ with the highest observed performance at $\lambda=80$ (Figure 6-B). The mean variance of pixels in the FMNIST training dataset is $\sigma_{\operatorname{input}}=0.087$. Although additive gaussian noise caused progressive deterioration in model performance from $\sigma_{\operatorname{noise}}=0.1$ to $\sigma_{\operatorname{noise}}=1$, test accuracy remained above $0.819$ for all $\sigma_{\operatorname{noise}}<0.4$, which corresponds to a signal-to-noise ratio of approximately $1:4.6$ (Figure 6-C). The impact of the randomised projections on model performance was assessed in 3-layer MLP models with $2\times 1000$ hidden neurons in the first two layers and $m_{3}\in\{250,500,1000\}$ in the penultimate layer, with stable test performance observed throughout (Figure 6-D).
Figure 6: A: Test accuracy of hidden layer explanations in a $4\times 1000$ neuron MLP trained by FP on FMNIST. B: Test accuracy of $3\times 1000$ neuron MLP trained trained by FP on FMNIST using different $\lambda$ regularisation values. C: Test accuracy of $3\times 1000$ neuron MLP trained trained by FP on FMNIST with additive Gaussian noise in training data. D: Variability of FP training performance using different random projections in $3\times 1000$ neuron MLP on FMNIST. Points represent mean over all replicates and error bars represent $\pm$ one standard deviation. BP: backpropagation; DTP: Difference Target Propagation; $F$: Forward-Forward; FP: Forward Projection; LP: label projection; LPN: noisy label projection; LS: Local Supervision; PC: Predicitve Coding; RF: random features.
A.9 Optimisation Theory
We now compare the fit of the Forward Projection estimator in a two-layer network, to that of a simple linear estimator in random data. We consider a dataset $\{\textbf{x}_{i},\textbf{y}_{i}\}_{i=1}^{N}$ with random Gaussian distribution such that $\textbf{x}_{i}\in\mathcal{N}(\textbf{0},\textbf{I}_{m})$ and an independent random normal target $\textbf{y}_{i}\in\mathcal{N}(\textbf{0},\textbf{I}_{d})$. We assume $N>m>d$. We consider a fully connected neural network with a single hidden layer of dimension $m$, with relu-activation. The task is to identify network weights $\{\textbf{W}_{1},\textbf{W}_{2}\}$ to minimise the error $\lVert\textbf{Y}-\operatorname{ReLU}(\textbf{XW}_{1})\textbf{W}_{2}\rVert$. We first consider the expected error of a single layer, linear model. For convenience, we will set the penalty factor $\lambda=0$, such that
$\textbf{W}_{1}:=(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}\textbf{Y}.$ (31)
We denote the projection matrix of X as
$\textbf{P}_{\textbf{x}}=\textbf{X}(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}.$ (32)
Therefore, the spectrum of $\textbf{P}_{\textbf{x}}$ is composed of a unit eigenvalue with multiplicity $m$ and a zero eigenvalue with multiplicity NmN-m. Letting $\hat{\textbf{Y}}_{1}=\textbf{P}_{\textbf{x}}\textbf{Y}$, we measure the error $\textbf{E}_{1}=\textbf{Y}-\textbf{P}_{\textbf{x}}\textbf{Y}$, such that
$\lVert\textbf{E}_{1}\rVert=\lVert(\textbf{I}-\textbf{P}_{\textbf{x}})\textbf{Y}\rVert.$ (33)
The idempotent matrix $(\textbf{I}-\textbf{P}_{\textbf{x}})$ has rank NmN-m, as it a projection matrix onto the orthogonal complement of the column space of X. Therefore, the expected squared Frobenius norm of the error is given by
$\mathbb{E}\left[\frac{\lVert\textbf{E}_{1}\rVert^{2}_{F}}{\lVert\textbf{Y}\rVert^{2}_{F}}\right]=\frac{N-m}{N}.$ (34)
Let $\textbf{U}\in\mathbb{R}^{d\times m}$ and $\textbf{Q}\in\mathbb{R}^{m\times m}$ . Letting $\tilde{\textbf{Z}}=\textbf{XQ}+\textbf{YU}$, we now fit $\textbf{W}_{1}$ to estimate $\tilde{\textbf{Z}}$ from X, such that
$\textbf{W}_{1}:=(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}(\textbf{XQ}+\textbf{YU}).$ (35)
As $(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}\textbf{X}=\textbf{I}$, we have
$\textbf{W}_{1}=\textbf{Q}+(\textbf{X}^{T}\textbf{X})^{-1}\textbf{X}^{T}\textbf{YU}.$ (36)
Thus, the pre-activation potential at the first layer is given by:
$\textbf{X}\textbf{W}_{1}=\textbf{XQ}+\textbf{P}_{\textbf{x}}\textbf{YU}.$ (37)
Assuming U has full row rank, we have $\textbf{UU}^{+}=\textbf{I}_{d}$, as $m>d$. Furthermore, As Q and U have random Gaussian entries, we have
$\mathbb{E}[\textbf{X}\textbf{W}_{1}\textbf{U}^{+}]=\mathbb{E}[\textbf{XQU}^{+}]+\mathbb{E}[\textbf{P}_{\textbf{x}}\textbf{YUU}^{+}]=\textbf{P}_{\textbf{x}}\textbf{Y}=\frac{1}{2}\hat{\textbf{Y}}_{1}.$ (38)
We now consider the covariance $\operatorname{Cov}(\textbf{XW}_{1},\textbf{Y})$. We assume $\textbf{XW}_{1}$ has zero mean. We have
$\mathbb{E}[\langle\textbf{XW}_{1},\textbf{Y}\rangle]=\mathbb{E}[\operatorname{tr}\left((\textbf{XQ}+\textbf{P}_{\textbf{x}}\textbf{YU})^{T}\textbf{Y}\right)]$ (39)
By independence of Q, we have $\mathbb{E}[\operatorname{tr}((\textbf{XQ})^{T}\textbf{Y})]=0$. The second term expands to
$\operatorname{tr}\left((\textbf{P}_{\textbf{x}}\textbf{YU})^{T}\textbf{Y}\right)=\operatorname{tr}\left(\textbf{U}^{T}\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y})\right)$ (40)
As $\textbf{P}_{\textbf{x}}$ is a projector matrix with rank $m$, and Y is Gaussian, we have
$\mathbb{E}[\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y}]=d\cdot\operatorname{tr}(\textbf{P}_{\textbf{x}})\textbf{I}_{d}=dm\cdot\textbf{I}_{d}$ (41)
Therefore
$\mathbb{E}[\langle\textbf{XW}_{1},\textbf{Y}\rangle]=\operatorname{tr}\left(\textbf{U}^{T}\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y})\right)=dm\cdot\operatorname{tr}(\textbf{U}^{T})$ (42)
As $m>d$ and U has random Gaussian entries, we have $\operatorname{Var}[\operatorname{tr}(\textbf{U}^{T})]=d$. Therefore, we have
$\mathbb{E}[\lvert\langle\textbf{XW}_{1},\textbf{Y}\rangle\rvert]=\lvert\operatorname{tr}\left(\textbf{U}^{T}\textbf{Y}^{T}\textbf{P}_{\textbf{x}}\textbf{Y})\right)\rvert\sim dm\sqrt{d}.$ (43)
After normalising for $\lVert\textbf{XW}_{1}\rVert_{F}\lVert\textbf{Y}\rVert_{F}\sim\sqrt{Nm}\cdot\sqrt{Nd}$, the correlation magnitude scales according to
$\mathbb{E}\left[\lvert\operatorname{Corr}(\textbf{XW}_{1},\textbf{Y})\rvert\right]\sim\frac{\operatorname{tr}\langle\textbf{XW}_{1},\textbf{Y}\rangle}{\lVert\textbf{XW}_{1}\rVert_{F}\lVert\textbf{Y}\rVert_{F}}=\frac{d\sqrt{m}}{N}.$ (44)
Assuming $\operatorname{rank}(\textbf{A})=m$, we may define $\textbf{P}_{\textbf{a}}=\textbf{A}(\textbf{A}^{T}\textbf{A})^{-1}\textbf{A}^{T}$. We note that A is dependent on Y with a non-linear relationship, which is induced by the $\operatorname{ReLU}$ activations. We now consider modelling Y in $\mathcal{C}(\textbf{A})\cap\mathcal{C}(\textbf{X})^{\perp}$, the subspace of A orthogonal to X. For large $N$, application of $\operatorname{ReLU}$ nonlinearities in the construction of A may yield $\operatorname{dim}\left(\mathcal{C}(\textbf{A})\cap\mathcal{C}(\textbf{X})^{\perp}\right)\approx m$. Recall from (44) that the column space of $\textbf{XW}_{1}$ is correlated with Y. As $\textbf{A}=\operatorname{ReLU}(\textbf{XW}_{1})$, the column space of A remains correlated with Y. Therefore
$\mathbb{E}\left[\frac{\lVert\textbf{Y}-\textbf{P}_{\textbf{a}}\textbf{Y}\rVert_{F}^{2}}{\lVert\textbf{Y}\rVert_{F}^{2}}\right]<\frac{N-m}{N}$ (45)
Therefore, even when modelling pure Gaussian noise, the error of the two-layer model is lower than that of the simple linear predictor.
A.10 Modelling high dimensional labels
From (28), we have:
For simplicity, we analyse the special case that $g_{l}$ is the identity function yielding,
$\hat{\textbf{y}}_{l}:=\left(\textbf{z}_{l}-\textbf{a}_{l-1}\textbf{Q}_{l}\right)\textbf{U}_{l}^{+}.$ (47)
Recalling that the realised potentials are estimates of pre-defined target potentials, such that $\textbf{z}_{l}=\tilde{\textbf{z}}_{l}-\textbf{e}_{l}$, we have
$\hat{\textbf{y}}_{l}:=\left(\tilde{\textbf{z}}_{l}-\textbf{e}_{l}-\textbf{a}_{l-1}\textbf{Q}_{l}\right)\textbf{U}_{l}^{+}.$ (48)
Applying the identity function as $g_{l}$ in equation (4), we have
$\tilde{\textbf{z}}_{l}=\textbf{a}_{l-1}\textbf{Q}_{l}+\textbf{y}\textbf{U}_{l}.$ (49)
Substituting, we have
$\hat{\textbf{y}}_{l}=\left(\textbf{y}\textbf{U}_{l}-\textbf{e}_{l}\right)\textbf{U}_{l}^{+}.$ (50)
Assuming, that U has full row rank, and that $m_{l}>m_{L}$, we have $\textbf{U}_{l}\textbf{U}^{+}_{l}=\textbf{I}_{m_{L}}$. Therefore
$\hat{\textbf{y}}_{l}=\textbf{y}-\textbf{e}_{l}\textbf{U}_{l}^{+}.$ (51)
Assuming $\textbf{e}_{l}\sim\mathcal{N}(0,\epsilon\textbf{I}_{m_{l}})$ for some $\epsilon\in\mathbb{R}$, we have $\mathbb{E}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]=0$, as U is random and independent of $\textbf{e}_{l}$. Therefore
$\mathbb{E}\left[\hat{\textbf{y}}_{l}\right]=\textbf{y}-\mathbb{E}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]=\textbf{y}.$ (52)
Therefore $\operatorname{Var}\left[\tilde{\textbf{y}}_{l}-\textbf{y}_{l}\right]=\operatorname{Var}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]$. Using the property of linear transformation of Gaussian variables, we have
$\operatorname{Var}\left[\textbf{e}_{l}\textbf{U}_{l}^{+}\right]=\epsilon\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}.$ (53)
We note that
$\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)=\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}$ (54)
Furthermore, it is noted that
$\mathbb{E}\left[\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}\right]=\mathbb{E}\left[\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)\right]\approx\mathbb{E}[\operatorname{Tr}\left((\textbf{U}_{l}\textbf{U}_{l}^{T})^{-1}\right)]$ (55)
We observe that the Frobenius norm $\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}$ equals the sum of inverse squared singular values
$\lVert\textbf{U}_{l}^{+}\rVert_{F}^{2}=\sum_{i=1}^{m_{L}}\frac{1}{\sigma_{i}}^{2}.$ (56)
Selecting $\textbf{U}_{l}$ with independent and identically distributed Gaussian entries, the singular values squared are the eigenvalues of the Wishart matrix
$\textbf{U}_{l}\textbf{U}_{l}^{T}\sim\mathcal{W}_{m_{L}}(m_{l},\textbf{I}_{m_{l}}).$ (57)
Therefore we have $\mathbb{E}[\operatorname{Tr}\left((\textbf{U}_{l}\textbf{U}_{l}^{T})^{-1}\right)]=\frac{m_{L}}{m_{l}-m_{L}-1}$ for $m_{L}<m_{l}$. Thus, the expectation $\mathbb{E}\left[\operatorname{Tr}\left(\textbf{U}_{l}^{+}\left(\textbf{U}_{l}^{+}\right)^{T}\right)\right]$ increases with $m_{L}$ and decreases with $m_{l}$. Thus, higher dimensional labels are subject to poorer encoding, for a given hidden layer dimension. Conversely, for a given label dimension, it is expected that representation and learning will improve with increases of the hidden layer dimension. Accordingly, adequate modelling of high dimensional labels may require larger hidden layer dimensions.
A.11 Modelling with Challenging Hidden Activation Functions
@@
Table 3: Test accuracies of local learning methods in FashionMNIST, Promoters and PTBXL-MI tasks, using modulo 2 (“mod2”) ($f(x)=x\mod{2}$) and square ($f(x)=x^{2}$) activation functions. FP: Forward Projection; LS: Local Supervision; $F$: Forward-Forward; BP: Backpropagation
A.12 Few-shot modelling
Discrimination performance of Difference Target Propagation and Predictive Coding methods in few-shot learning tasks is plotted in Figure 7. Both methods proved uninformative for few-shot learning CXR (Figure 7-A), OCT (Figure 7-B), and CIFAR tasks (Figure 7-C).
Figure 7: Test performance of few-shot trained 2D-CNN models, showing Difference Target Propagation (DTP) and Predictive Coding (PC) methods. Mean test AUC is reported over 50 few-shot training experiments. A: Chest X-ray (CXR) task. B: Optical Coherence Tomography (OCT) task. C: CIFAR task in which models were required to classify the first two classes (aeroplane and automobile). Mean test AUC is reported over 50 few-shot training experiments. Models were fitted with $N\in\{5,10,15,20,30,40,50\}$ training samples from each class in CXR and OCT tasks and and $N\in\{25,50,75,100\}$ samples per class for the CIFAR task.
@@ A.12 Few-shot modelling
Discrimination performance (AUC) of Forward Projection, random features, Local Supervision, Forward-Forward, Predictive Coding, Difference Target Propagation and and backpropagation in few-shot modelling tasks is provided in Supplementary Table 4.
AUC
Method FP (ours) RF LS $F$ PC DTP BP
Dataset Partition $N_{\text{train}}$
CXR Train N=5 100.0 ± 0.0 100.0 ± 0.0 50.7 ± 16.0 52.0 ± 9.5 51.0 ± 12.9 53.9 ± 12.7 72.5 ± 9.4
N=10 100.0 ± 0.0 100.0 ± 0.0 48.7 ± 12.7 48.8 ± 6.3 49.4 ± 9.9 53.8 ± 10.4 69.0 ± 8.1
N=15 100.0 ± 0.0 100.0 ± 0.0 55.2 ± 13.9 51.1 ± 6.1 49.4 ± 9.5 50.4 ± 8.8 72.4 ± 10.6
N=20 100.0 ± 0.0 100.0 ± 0.2 58.5 ± 12.8 50.0 ± 5.5 50.2 ± 9.6 52.0 ± 7.4 72.6 ± 8.0
N=30 100.0 ± 0.0 99.9 ± 0.4 59.8 ± 15.5 50.0 ± 5.7 52.4 ± 9.0 49.4 ± 8.4 71.8 ± 9.6
N=40 100.0 ± 0.1 99.3 ± 0.8 61.4 ± 15.8 49.7 ± 4.8 52.1 ± 8.9 52.3 ± 7.9 74.3 ± 10.2
N=50 99.9 ± 0.2 98.4 ± 1.0 66.1 ± 14.8 49.4 ± 6.2 51.4 ± 8.2 52.2 ± 7.3 84.2 ± 9.8
Test N=5 75.2 ± 7.0 67.6 ± 7.5 49.7 ± 8.8 49.3 ± 5.7 50.2 ± 8.1 51.8 ± 7.6 57.6 ± 7.6
N=10 76.6 ± 5.8 68.1 ± 6.0 47.9 ± 9.0 49.1 ± 5.7 49.3 ± 8.9 52.7 ± 8.3 59.7 ± 6.2
N=15 77.4 ± 4.5 69.2 ± 4.8 53.9 ± 12.3 49.9 ± 6.7 49.0 ± 8.3 50.2 ± 7.6 63.1 ± 8.6
N=20 78.9 ± 3.7 69.0 ± 4.9 58.9 ± 11.5 49.0 ± 5.8 49.6 ± 8.5 51.0 ± 7.1 66.1 ± 9.3
N=30 80.6 ± 3.7 69.5 ± 3.7 59.2 ± 14.0 51.2 ± 6.2 51.9 ± 8.8 49.1 ± 8.9 66.6 ± 8.0
N=40 81.7 ± 2.9 72.4 ± 3.7 61.2 ± 16.1 48.0 ± 6.4 50.3 ± 7.6 51.3 ± 8.6 70.2 ± 8.3
N=50 82.5 ± 3.5 72.2 ± 3.1 64.8 ± 14.9 49.4 ± 7.4 50.1 ± 7.9 50.9 ± 7.2 78.9 ± 8.3
OCT Train N=5 100.0 ± 0.0 100.0 ± 0.0 53.9 ± 27.2 49.4 ± 12.5 52.5 ± 23.0 56.9 ± 21.5 86.8 ± 13.2
N=10 100.0 ± 0.0 100.0 ± 0.0 58.3 ± 17.8 50.4 ± 12.2 51.4 ± 15.9 54.4 ± 13.8 83.3 ± 12.2
N=15 100.0 ± 0.0 100.0 ± 0.0 57.1 ± 15.7 49.9 ± 11.2 49.6 ± 12.3 52.5 ± 13.8 80.5 ± 11.2
N=20 100.0 ± 0.0 100.0 ± 0.0 53.4 ± 21.0 51.6 ± 9.8 52.9 ± 13.0 52.8 ± 11.7 88.9 ± 12.1
N=30 100.0 ± 0.0 100.0 ± 0.0 58.0 ± 18.8 50.1 ± 11.3 50.1 ± 11.5 51.7 ± 9.7 83.7 ± 12.3
N=40 100.0 ± 0.0 99.9 ± 0.3 61.4 ± 21.0 52.3 ± 9.2 52.2 ± 11.2 51.0 ± 12.6 88.0 ± 12.0
N=50 100.0 ± 0.0 99.7 ± 0.7 61.7 ± 19.8 50.8 ± 13.6 51.1 ± 8.8 53.6 ± 11.0 95.2 ± 8.3
Test N=5 76.3 ± 11.5 63.1 ± 10.0 51.1 ± 17.6 52.5 ± 12.0 50.1 ± 15.6 47.1 ± 10.5 61.8 ± 17.4
N=10 84.5 ± 5.8 66.5 ± 7.7 54.6 ± 16.7 51.8 ± 12.1 49.0 ± 17.1 49.2 ± 11.6 71.8 ± 15.9
N=15 86.5 ± 7.8 68.7 ± 7.4 56.8 ± 16.3 47.8 ± 12.8 49.6 ± 13.6 52.2 ± 12.9 72.3 ± 14.4
N=20 88.4 ± 5.9 69.5 ± 6.0 49.6 ± 19.1 52.3 ± 13.0 50.2 ± 15.5 49.9 ± 14.7 82.4 ± 14.4
N=30 89.9 ± 4.0 70.5 ± 5.7 55.7 ± 20.9 51.8 ± 11.7 49.6 ± 15.5 48.4 ± 13.9 81.9 ± 12.1
N=40 90.2 ± 5.0 73.3 ± 5.5 62.0 ± 25.1 50.6 ± 13.0 51.6 ± 16.5 51.1 ± 15.4 88.7 ± 9.3
N=50 91.2 ± 4.7 73.6 ± 5.9 62.6 ± 23.8 49.7 ± 18.4 51.2 ± 12.9 51.2 ± 13.4 94.2 ± 6.6
CIFAR Train N=25 95.3 ± 2.7 100.0 ± 0.0 71.6 ± 23.6 52.6 ± 9.7 49.6 ± 17.5 55.9 ± 16.2 96.8 ± 4.5
N=50 96.0 ± 2.0 100.0 ± 0.0 78.9 ± 18.3 59.2 ± 21.0 60.3 ± 16.5 56.1 ± 13.8 99.4 ± 1.9
N=75 96.0 ± 1.6 100.0 ± 0.0 86.7 ± 5.2 67.5 ± 17.8 64.6 ± 15.1 63.4 ± 11.4 98.2 ± 2.5
N=100 95.9 ± 1.3 100.0 ± 0.1 88.7 ± 3.8 73.7 ± 10.6 69.6 ± 11.0 60.9 ± 8.9 99.9 ± 0.3
Test N=25 78.6 ± 3.4 78.6 ± 3.6 65.0 ± 20.5 52.5 ± 8.8 47.7 ± 14.6 54.2 ± 14.1 82.2 ± 2.4
N=50 83.1 ± 2.8 77.3 ± 3.1 74.2 ± 16.5 60.1 ± 22.8 58.5 ± 16.3 54.6 ± 13.4 86.3 ± 2.2
N=75 84.4 ± 1.7 76.2 ± 3.1 81.5 ± 2.6 68.3 ± 18.8 64.4 ± 15.0 62.8 ± 10.7 87.5 ± 2.1
N=100 86.1 ± 2.0 78.1 ± 2.6 83.5 ± 2.8 74.3 ± 9.7 68.6 ± 11.9 60.7 ± 9.4 89.5 ± 1.4
@@
Table 4: Train and Test AUC performance in few-shot learning experiments. AUC: area under curve. CXR: chest x-ray dataset. OCT: optical coherence tomography dataset. BP: Backpropagation; DTP: Difference Target Propagation; $F$: Forward Forward; FP: Forward Projection; LS: Local Supervision; PC: Predictive Coding; RF: Random features
A.13 Training Times
Training times and epochs for each method, in full dataset modelling. FMNIST was modelled with a $3\times 1000$ neuron MLP network. Sequential datasets (PTBXL-MI and Promoters) were modelled by a 1d-CNN architecture of four convolutional blocks with $32\times 2^{l-1}$ filters in the $l$-th block. CIFAR modelling used a vision transformer architecture operating on image patches of dimension $4\times 4$, with a sequential stack of four multi-headed attention layers, each having 8 heads, embedding dimension 64, and MLP dimension 64. All experiments were run on the Google Colab service using an NVIDIA T4 graphics processing unit.
Method FP RF LS $F$ PC DTP BP
Dataset Metric
FMNIST Time (s) 0.3 ± 0.1 0.1 ± 0.0 46.9 ± 12.5 171.9 ± 56.6 150.2 ± 61.2 72.7 ± 26.4 22.4 ± 4.8
Promoters 6.5 ± 0.2 3.3 ± 0.6 91.6 ± 37.7 69.5 ± 12.4 258.9 ± 233.7 103.0 ± 23.6 30.9 ± 4.3
PTBXL-MI 10.6 ± 1.1 6.7 ± 0.8 56.3 ± 13.8 35.5 ± 11.9 73.3 ± 15.8 61.0 ± 24.2 18.7 ± 5.1
CIFAR 16.3 ± 0.6 9.2 ± 0.6 43.3 ± 15.1
FMNIST Epochs 1.0 ± 0.0 1.0 ± 0.0 10.8 ± 3.6 40.2 ± 13.8 25.4 ± 10.2 13.6 ± 5.3 11.6 ± 2.7
Promoters 1.0 ± 0.0 1.0 ± 0.0 11.4 ± 4.9 6.2 ± 1.3 18.4 ± 17.5 7.8 ± 2.0 10.0 ± 1.4
PTBXL-MI 1.0 ± 0.0 1.0 ± 0.0 17.0 ± 4.3 7.2 ± 2.3 14.6 ± 3.3 13.8 ± 5.9 13.2 ± 3.6
CIFAR 1.0 ± 0.0 1.0 ± 0.0 10.8 ± 3.9
Table 5: Training times and epochs in full dataset modelling experiments. BP: Backpropagation; DTP: Difference Target Propagation; $F$: Forward Forward; FP: Forward Projection; LS: Local Supervision; PC: Predictive Coding; RF: Random features